OpenStudy (anonymous):

Center City East Parking Garage has a capacity of 258 cars more than Center City West Parking Garage. If the combined capacity for the two garages is 1224 cars, find the capacity for each garage.

OpenStudy (anonymous):

You need to know how to do these and not just the answers, right? :) times=multiply sum=add -=subtract 6 (x-150)=36 divide both sides by 6 x-150=6 add 150 to both sides x=6+150 x=156 6(x-150)=36 6(156-150)=36 6(6)=36 2 Mary has 3 times John mary=3j John=j Mary plus John=160 3j+j=160 4j=160 divide both sides by 4 j=40 3) number-22=5 solve this one like you've done the others! 4) 3 times (n+3)=5times n-9 This should get you going :)

OpenStudy (anonymous):

w + ( w + 258 ) = 1224 cars -> 2w = 1224 - 258 => w = 966/2 = 483 cars in West parking Thus East parking capacity is: 483 + 258 = 741 cars